OpenStudy (alibaby):

Simplify 10^6 over 10^-3 one divided by the quantity ten to the third one divided by the quantity ten to the eighteenth 103 109

OpenStudy (3mar):

\[10^{9}\]

rishavraj (rishavraj):

\[\frac{ 10^6 }{ 10^{-3} } = 10^{6 - (- 3)} = 10^{6 + 3} = 10^9\]

OpenStudy (3mar):

Thanks for details, rishavraj !